### v2

@@ 2.3.1 AdaPRL for multi-task learning
Obviously, there are several problems with $L_{MCPRL_{nav}}$. The loop over batch and targets will lead to a significant increase in computational complexity and memory usage, especially when $N$ is relatively large. More importantly, the data type, endogenous meaning, as well as magnitude between targets could be vastly different, therefore not suitable for calculating differences.
We propose an improved version of the multi-output version of $L_{CPRL}$. Instead of considering across all targets over every samples, we only focus on pairs within each target. Assume the number of target is $N$, the elements $(i,j)$ of $kth$ matrix $M$ corresponding to whether sample $i$ and sample $j$ are a valid ranked pair on target $k$, and elements $(i,j)$ of $kth$ matrix $C$ indicate the level of confidence in the certainty of the sample pair $i$ and $j$ on target $k$. Formally, we obtain:
@@
Since only the sample pairs within each target are considered, it is evident that the computational complexity reduces from $O\left((BN)^{2}\right)$ to $O\left(B^{2}N\right)$. Therefore, the approach not only ensures that comparisons of data pairs are meaningful but also significantly improves computation efficiency.
@@ 2.3.2 AdaPRL for multivariate time series forecasting
Time series forecasting is a crucial task in various domains, and deep learning models have significantly advanced research in this area. Multivariate time series forecasting is a major subtopic in time series forecasting. Multivariate time series forecasting involves predicting future values of multiple interrelated time-dependent variables simultaneously, emphasizes the importance to analyze and model the relationships between several targets that change over time. Similar to our analysis of multi-task learning before, to integrate AdaPRL in multivariate time series forecasting problem, the naive method of comparing all targets over all prediction length in every possible pairs is both computationally expensive and memory-intensive. Instead, we focus on the inner correlation of every target over time dimension. That is to say:
@@
The computational complexity reduces from $O\left((BNT)^{2}\right)$ to $O\left((BT)^{2}N\right)$ comparing to naive pairwise learning applied to multivariate time series forecasting. Since the number of targets in many multivariate datasets could be relatively large, this adoption could save computing and memory and work well with short-term time series forecasting.
However, in the context of long-term time series forecasting, the additional computational complexity remains significant because the prediction sequence could be $720$ or even longerOreshkin et al. (2019); Nie et al. (2022); Zhang & Yan (2023); Zeng et al. (2023). To address this challenge, we further propose a novel sparse confidence pairwise learning, annotated as $L_{SCPRL}$. The definition is:
@@
where $S$ indicates a sparse area that can be derived either randomly or heuristically. Note that the area of $S$ satisfies $|S|\ll B^{2}T^{2}$, therefore, the computational complexity and memory consumption reduce significantly from $O\left((BT)^{2}N\right)$ to $O\left(|S|N\right)$.
Furthermore, if randomness is induced when generating the sparse area $S$, $L_{SCPRL}$ can also be viewed as a variant of dropout Srivastava et al. (2014) in the Confidence Pairwise Regression Loss (CPRL). Conclusively, our sparse confidence pairwise learning approach only conducts comparison within certain area instead of all pairs, takes most advantage of learning patterns over a longer time horizon while theoretically reducing overfitting and remaining computation and memory efficiency. Note that sparse confidence pairwise learning loss can be integrated with not just multivariate time series forecasting frameworks, but also single-output models, multi-task learning models and other sophisticated models simply by adjusting the definition of sparse area $S$ for the purpose of reducing overfitting.
@@
Table 1: Performance comparison of all methods across regression datasets concerning Mean Squared Error (MSE), Mean Absolute Error (MAE), and Kendall’s $\tau$. We bolded the best performance, while underlined scores were the second best of each dataset. Rel.Imp indicated the relative improvement of AdaPRL over the baseline. The results of RankSim and Ordinal Entropy for the ETT and Janestreet datasets were not reported because they were not compatible with the time series forecasting task. Detailed performance of each method on individual datasets was elaborated in the experiment section of the Appendix.
@@ 3 Experiments
In this section, we conducted extensive experiments on a variety of regression datasets in different fields for the purpose of demonstrating that this method had universal generalization capabilities not only across different datasets but also across different backbone networks. Datasets included tabular regression datasets, natural language, and vision datasets. The research field of data also varied from recommendation, time-series prediction, age prediction, natural language understanding, finance, etc. Moreover, we conducted offline evaluations and online AB tests with millions of users, which proved that our method also benefited real-world industrial applications. Due to space limitations, we presented only the overall results for each dataset in the main body of the paper. Comprehensive experimental details were thoroughly discussed in the Appendix.
@@ 3.1 Experiment setup
To comprehensively evaluate the proposed method, an extensive experiment was conducted to assess the performance across multiple real-world datasets. The data domains included recommendation systems, time series forecasting, age prediction, natural language understanding, finance, and industrial data. Specifically, the evaluation datasets included Movielens-1M, KuaiRec, ETT, AgeDB, CLAP2016, Jigsaw, Janestreet, and a large-scale proprietary industrial dataset from a major music streaming app. The statistics of the datasets are summarized in Table 2. We choose MSE loss as baseline method, denoted as Baseline(L2), and we choose RankSimGong et al. (2022) and Ordinal EntropyZhang et al. (2023) for comparison because they are claimed to be able to effective in regression tasks. We also adapt the original form of pairwise regression loss without uncertainty estimation described in equations 4 and 5 for comparison. Please refer to the Appendix for detailed experiment results and experiment setups.
@@
Time series forecast. ETT(4 subsets) 174,204 7 5,040
@@
Table 2: Statistics of datasets used in the experiments section. Note that #Features means differently in different domain’s dataset. For recommendation data it indicates the number of sparse features, while for image datasets and nature language data, it indicates the number of pixels and number of tokens of each sample respectively.
@@ 3.2 Overall Performance
As shown in Table 1, we compared several commonly used regression loss functions on various datasets of multiple domains, including the Mean Squared Error (MSE) (also known as the $L2$) loss, as well as regression learning methods such as RankSimGong et al. (2022) and Ordinal EntropyZhang et al. (2023) method, alongside our proposed methods pairwise regression learning (PRL) and AdaPRL. We conducted a comprehensive horizontal comparison of the key performance of these regression methods across various datasets, including Movielens-1M, KuaiRec, ETT, ECL, AgeDB, CLAP2016, Jigsaw, Janestreet, and industrial dataset. As indicated in the table, our proposed method AdaPRL achieved the best performance on the vast majority of datasets, surpassing the previously introduced regression learning methods RankSim and ordinal entropy in all datasets except KuaiRec, thereby demonstrating significant potential and generalization capabilities. In addition, the pairwise regression learning (PRL) method also exhibited a certain degree of improvement over the original MSE baseline on most datasets. Furthermore, Kendall’s $\tau$ of our proposed is also higher compared with traditional loss functions, indicating that it further improves ranking ability which helps distinguish the monotonicity of data.
@@ 3.3 Online A/B testing
To evaluate the effectiveness of the AdaPRL algorithm in real-world commercial applications, we deployed it with tuned hyperparameters in a music streaming application that serves tens of millions of active users every day. The regression task is a crucial component of the value estimation module within our user-targeting framework, which is integral to a commercial business generating millions in daily revenue, therefore even a minor improvements in estimation accuracy can yield significant commercial benefits. Specifically, we conducted an online A/B test comprising a control group using the original mean squared error loss and an experimental group applying the AdaPRL algorithm in the regression component, each representing 10% of our total traffic of the APP. Over a two-week period, we measured the actual revenue per user generated by both groups. As shown in Table 3, the average daily income per user resulting from the use of the AdaPRL algorithm in regression increased by 2.57% compared to the same model using the MSE loss, and the increase is statistically significant according to the t-test with p-value $<$ 0.05, suggesting that AdaPRL can enhance commercial benefits in real world applications.
@@ 3.4 Further Study and Analysis
To facilitate a deeper understanding of the underlying mechanisms, such as the sensitivity of hyperparameter,the effect of sparsity, we conducted several comprehensive experiments for our proposed method. Additionally, various experimental analyzes were performed to examine the model’s robustness and generalizability across different scenarios and datasets, which demonstrates that AdaPRL improves the model’s robustness to label noise of training data, resilience to unforeseen data corruption and the robustness to reduced training data. Moreover, extensive experiments show that AdaPRL can be seamlessly incorporated into various regression frameworks, such as RankSim and Ordinal Entropy, and boost their performance. In addition, we also used the visualization diagram on the time series forecasting task to show that AdaPRL not only improves the accuracy but also obtains the confidence interval of the prediction through uncertainty estimation, thereby improving the interpretability of the prediction results.
Due to the limited page, the detailed experimental settings,results, discussions and further study are relegated to the Appendix.
@@ 4 Conclusion
In this paper, we propose a novel pairwise regression learning framework called AdaPRL which leverages the relative differences between data points and integrates with deep probabilistic models to quantify the uncertainty associated with the predictions. Extensive experiments, as well as online A/B testing in a major music streaming APP, demonstrate that our method has the potential to enhance the accuracy, robustness, and interpretability of modern deep learning regression models across various domains, including recommendation systems, time series forecasting, age prediction, natural language processing, financial, and real-world industrial applications.
@@
This paper focuses on proposing a robust and generalized pairwise learning algorithm to enhance the performance of regression tasks in various domains of deep learning and push the accuracy limit to a higher level in all types of regression tasks with different backbones. We believe that this work is helpful to advance this goal. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1.1 Datasets
MovieLens-1M [^1] https://grouplens.org/datasets/movielens/1m/Harper & Konstan (2015) is a popular dataset in recommendation system research that contains 1 million records of the (user-features, item-features, rating) triplet, where the rating scores are between 1 and 5. Similar to previous work Song et al. (2019); Wang et al. (2020), we also formalize the task as a rating regression problem. We use all single-value categorical features as the input of embedding layer, followed by a multilayer perceptron to predict the target. The data is randomly split into 80% for training, 10% for validation and 10% for testing.
KuaiRec [^2] https://kuairec.com/Gao et al. (2022) is a real-world dataset collected from the recommendation logs of a famous mobile video sharing app with fully observed user-item interaction matrix that originally aims to be used in unbiased offline evaluation for recommendation models. We employ the fully observed small matrix for both training, validation, and testing. All user and item features are categorized using feature binning. These features are then processed through an embedding layer, followed by a multilayer perceptron, to predict video completion rates. The data is randomly split into 80% for training, 10% for validation, and 10% for testing.
@@ A.1.2 Experiment Setup
We performed early stop strategy with a patience of three in the valid set using MSE metrics. Every experiment was repeated five times and the average of corresponding metrics in terms of mean squared error (MSE), mean absolute error (MAE), and Kendall’s $\tau$ was reported. A coarse-grained grid search was performed on the hyperparameters, including pairwise ranking loss type (MAE or RMSE type as described in 3.2.2) and the weight factor $\alpha$ between {0.01,0.02,0.05,0.1,0.2,0.5}.
@@
Table 4: Regression accuracies for Pairwise Regression Learning and AdaPRL on ML-1M. Rel.Imp indicated relative improvement over corresponding metrics. The experiments were repeated five times, and the average results were reported.
@@
Table 5: Regression accuracies for Pairwise Regression Learning and AdaPRL on KuaiRec. Rel.Imp indicated relative improvement over corresponding metrics. The experiments were repeated five times, and the average results were reported.
@@ A.1.3 Overall Performance
We categorized the dataset and bolded the best performance, while underlined scores were the second best. The experimental results were shown in Table 4 and 5, and we could draw the following conclusions: by comparing using single $L2$ loss, and $L2$ with original pairwise learning ($L2+PRL$) and AdaPRL method, the AdaPRL achieved highest accuracy all other methods in both Movielens-1M. Specifically, the test MSE and MAE metrics were reduced by 14.82% and 4.74%, 2.05% and 1.43% and 0.46% and 0.34% with respectively with $L1$, $Huber$ and $L2$ on the ML-1M dataset. AdaPRL also achieved best performance in KuaiRec dataset combining with three different $L_{reg}$, the test MSE and MAE metrics were reduced by 5.01% and 6.76%, 1.51% and 2.42% and 0.83% and 6.93% with respectively with $L1$, $Huber$ and $L2$ on the KuaiRec dataset. Moreover, the Kendall’s $\tau$ increased by 0.45% and 3.05% respectively, on the ML-1M and KuaiRec datasets with $L2$, indicating that the model not only generated more accurate predictions but also provided a stronger monotonic relationship between labels and predictions.
A.2 Time Series forecasting datasets

### v3

@@ 2.3.1 AdaPRL for multi-task learning
Obviously, there are several problems with $L_{MCPRL_{nav}}$. The loop over batch and targets will lead to a significant increase in computational complexity and memory usage, especially when $N$ is relatively large. More importantly, the data type, endogenous meaning, as well as magnitude between targets could be vastly different therefore not suitable for calculating differences.
To solve this issue, we propose an improved version of $L_{CPRL}$ for multi-output networks. Instead of considering across all targets over every sample, we only focus on pairs within each target space. Assume the number of tasks is $N$, we define a series of matrices $\{C\}^{k}$ and $\{M\}^{k}$. The elements $(i,j)$ of $k$th matrix $M$ corresponding to whether sample $i$ and sample $j$ are a valid ranked pair on target $k$, and elements $(i,j)$ of $k$th matrix $C$ indicate the level of confidence in the certainty of the sample pair $i$ and $j$ on target $k$. Formally, we obtain:
@@
Since only the sample pairs within each target are considered, it is evident that the computational complexity reduces from $O\left((BN)^{2}\right)$ to $O\left(B^{2}N\right)$. Therefore, the approach not only ensures that comparisons of data pairs are meaningful but also significantly improves computational efficiency.
@@ 2.3.2 AdaPRL for multivariate time series forecasting
Time series forecasting is a crucial task in various domains, and deep learning models have significantly advanced research in this area. Multivariate time series forecasting is a major subtopic in time series forecasting that involves predicting future values of multiple interrelated time-dependent variables simultaneously, emphasizing the importance of analyzing and modeling the relationships between several targets that change over time.
Similar to our analysis of multi-task learning above, to integrate AdaPRL in the multivariate time series forecasting, the naive method that compares all targets over all prediction lengths in every possible pair would be computationally expensive and memory-intensive. Instead, we focus on the inner correlation of every target over the time dimension:
@@
The computational complexity reduces from $O\left((BNT)^{2}\right)$ to $O\left((BT)^{2}N\right)$ compared to naive pairwise learning applied to multivariate time series forecasting. Since the number of targets in many multivariate datasets could be relatively large, this adoption could save computing and memory and work well with short-term time series forecasting.
However, in the context of long-term time series forecasting, the additional computational complexity remains significant because the prediction sequence could be $720$ or even longer Oreshkin et al. (2019); Nie et al. (2022); Zhang & Yan (2023); Zeng et al. (2023). To address this challenge, we further propose a novel Sparse Confidence Pairwise Learning (SCPRL), annotated as $L_{SCPRL}$. The definition is:
@@
where $S$ indicates a sparse area that can be derived either randomly or heuristically. Note that the area of $S$ satisfies $|S|\ll B^{2}T^{2}$; therefore, the computational complexity and memory consumption reduce significantly from $O\left((BT)^{2}N\right)$ to $O\left(|S|N\right)$.
Furthermore, if randomness is induced when generating the sparse area $S$, $L_{SCPRL}$ can also be viewed as a variant of dropout Srivastava et al. (2014) in Confidence Pairwise Regression Loss of 13. Conclusively, our sparse confidence pairwise learning approach only conducts comparisons within a certain area instead of all pairs, takes the most advantage of learning patterns over a longer time horizon while theoretically reducing overfitting and maintaining computation and memory efficiency. Note that sparse confidence pairwise learning loss can be integrated with not just multivariate time series forecasting frameworks, but also single-output, multi-task learning and other sophisticated models simply by adjusting the definition of sparse area $S$.
@@
Table 1: Performance comparison of all methods across regression datasets concerning Mean Squared Error (MSE), Mean Absolute Error (MAE), and Kendall’s $\tau$. We bold the best performance, while the underlined scores were the second best of each dataset. Rel.Imp indicates the relative improvement of AdaPRL over the baseline. The results of RankSim and Ordinal Entropy for the ETT and Janestreet datasets were not reported because they were not compatible with the time series forecasting task. Detailed performance of each method on individual datasets was elaborated in the experiment section of the Appendix.
@@ 3 Experiments
In this section, we conducted extensive experiments on a variety of regression datasets in different fields for the purpose of demonstrating that this method had universal generalization capabilities not only across different datasets but also across different backbone networks. Datasets included tabular regression datasets, natural language, and vision datasets. The research field of data also varied from recommendation, time-series prediction, age prediction, natural language understanding, finance, etc. Moreover, we conducted offline evaluations and online AB tests with millions of users, which proved that our method also benefited real-world industrial applications. Due to page limitations, we presented only the overall results for representative datasets in the main body of the paper. Comprehensive experimental details were thoroughly discussed in the Appendix A.
@@ 3.1 Experiment Setup
To comprehensively evaluate the proposed method, an extensive experiment was conducted to assess the performance across multiple real-world datasets. The data domains included recommendation systems, time series forecasting, age prediction, natural language understanding, finance, and industrial data. Specifically, the evaluation datasets included Movielens-1M, KuaiRec, ETT, AgeDB, CLAP2016, Jigsaw, Janestreet, and a large-scale proprietary industrial dataset from a major music streaming app. The statistics of the datasets are summarized in Table 2. We chose MSE loss as the baseline method, denoted as Baseline(L2), and we chose RankSim Gong et al. (2022) and Ordinal Entropy Zhang et al. (2023) for comparison because they were claimed to be effective in regression tasks. We also adapted the original form of pairwise regression loss without uncertainty estimation described in equations 4 and 5 for comparison. Please refer to the Appendix A for detailed experiment setups as well as the experiment results for all datasets.
@@
Time series forecast. ETT 174,204 7 5,040
@@
Table 2: Statistics of datasets used in the experiments section. Note that #Features means differently in different domain’s dataset. For recommendation data it indicates the number of sparse features, while for image datasets and nature language data, it indicates the number of pixels and number of tokens of each sample respectively, while #Fields denotes number of individual columns for tabular data and denotes number of task in time series forecasting dataset.
@@ 3.2 Overall Performance
As shown in Table 1, we compared several commonly used regression loss functions on various datasets of multiple domains, including the Mean Squared Error (MSE) (also known as the $L2$) loss, as well as regression learning methods such as RankSim Gong et al. (2022) and Ordinal Entropy Zhang et al. (2023) method, alongside our proposed methods pairwise regression learning (PRL) and AdaPRL. We conducted a comprehensive horizontal comparison of the key performance of these regression methods across various datasets, including Movielens-1M, KuaiRec, ETT, ECL, AgeDB, CLAP2016, Jigsaw, Janestreet, and an industrial dataset. As indicated in the table, our proposed method AdaPRL achieved the best performance on the vast majority of datasets, surpassing the previously introduced regression learning methods RankSim and Ordinal Entropy in all datasets except KuaiRec, thereby demonstrating significant potential and generalization capabilities. In addition, the pairwise regression learning (PRL) method also exhibited a certain degree of improvement over the original MSE baseline on most datasets. Furthermore, Kendall’s $\tau$ of our proposed method was also higher compared with traditional loss functions, indicating that it further improves ranking ability, which helps distinguish the monotonicity of data.
@@ 3.3 Online A/B Testing
To evaluate the effectiveness of the AdaPRL algorithm in real-world commercial applications, we deployed it with tuned hyperparameters in a music streaming application that serves tens of millions of active users every day. The regression task is a crucial component of the value estimation module within our user-targeting framework, which is integral to a commercial business generating millions in daily revenue; therefore, even minor improvements in the accuracy can yield significant commercial benefits. Specifically, we conducted an online A/B test comprising a control group using the original mean squared error loss and an experimental group applying the AdaPRL algorithm in the regression component, each representing 10% of our total traffic of the App. Over a two-week period, we measured the actual revenue per user generated by both groups. As shown in Table 3, the average daily income per user resulting from the use of the AdaPRL algorithm in regression increased by 2.57% compared to the same model using the MSE loss, and the increase was statistically significant according to the t-test with p-value $<$ 0.05, suggesting that AdaPRL could enhance commercial benefits in real-world applications.
@@ 3.4 Further Study and Analysis
To facilitate a deeper understanding of the underlying mechanisms, such as the sensitivity of hyperparameters,the effect of sparsity, we conducted several comprehensive experiments for our proposed method. Additionally, various experimental analyses were performed to examine the model’s robustness and generalizability across different scenarios and datasets, which demonstrates that AdaPRL improves the model’s robustness to label noise in the training data, resilience to unforeseen data corruption, and robustness to reduced training data. Moreover, extensive experiments show that AdaPRL can be seamlessly incorporated into various regression frameworks, such as RankSim and Ordinal Entropy, and boost their performance. In addition, we visualize the results on the time series forecasting task to demonstrate that AdaPRL not only enhances predictive accuracy but also provides confidence intervals for predictions through uncertainty quantification, thereby improving the interpretability of the results. Theoretical and experimental analysis shows that the extra computational cost of AdaPRL is generally considered acceptable. Due to the limited page, the detailed experimental settings, results, discussions, and further study are relegated to the Appendix.
@@ 4 Conclusion
In this paper, we propose a novel pairwise regression learning framework called AdaPRL, which leverages the relative differences between data points and integrates with deep probabilistic models to quantify the uncertainty associated with the predictions. Extensive experiments, as well as online A/B testing in a major music streaming App, demonstrate that our method has the potential to enhance the accuracy, robustness, and interpretability of modern deep learning regression models across various domains, including recommendation systems, time series forecasting, age prediction, natural language processing, finance, and real-world industrial applications.
@@
This paper focuses on proposing a generalized and robust pairwise learning algorithm to enhance the performance of universal regression tasks in various domains of deep learning, and push the accuracy limit to a higher level. We believe that this work is helpful to advance the goal. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1.1 Datasets
MovieLens-1M [^1] https://grouplens.org/datasets/movielens/1m/ Harper & Konstan (2015) is a popular dataset in recommendation systems research that contains 1 million records of the (user-features, item-features, rating) triplet, where the rating scores are between 1 and 5. Similar to previous work Song et al. (2019); Wang et al. (2020), we also formalize the task as a rating regression problem. We use all single-value categorical features as the input of the embedding layer, followed by a multilayer perceptron to predict the target. The data is randomly split into 80% for training, 10% for validation, and 10% for testing.
KuaiRec [^2] https://kuairec.com/ Gao et al. (2022) is a real-world dataset collected from the recommendation logs of a famous mobile video sharing app with a fully observed user-item interaction matrix that originally aims to be used in unbiased offline evaluation for recommendation models. We employ the fully observed small matrix for both training, validation, and testing. All user and item features are categorized using feature binning. These features are then processed through an embedding layer, followed by a multilayer perceptron, to predict video completion rates. The data is randomly split into 80% for training, 10% for validation, and 10% for testing.
@@ A.1.2 Experiment Setup
We performed an early stopping strategy with a patience of three in the valid set using MSE metrics. Every experiment was repeated five times and the average of corresponding metrics in terms of mean squared error (MSE), mean absolute error (MAE), and Kendall’s $\tau$ was reported. To demonstrate the robustness of our method to hyperparameters, we only performed a coarse-grained grid search in most of our experiments. The hyperparameters include pairwise ranking loss type (MAE or RMSE type as described in 3.2.2) and the weight factor $\alpha$ between {0.01,0.02,0.05,0.1,0.2,0.5}.
@@
Table 4: Regression accuracies for Pairwise Regression Learning and AdaPRL on ML-1M. Rel.Imp indicates relative improvement over corresponding metrics. The experiments were repeated five times, and the average results were reported.
@@
Table 5: Regression accuracies for Pairwise Regression Learning and AdaPRL on KuaiRec. Rel.Imp indicates relative improvement over corresponding metrics. The experiments were repeated five times, and the average results were reported.
@@ A.1.3 Overall Performance
We categorized the dataset and bolded the best performance, while the underlined scores were the second best. The experimental results are shown in Table 4 and 5. We can draw the following conclusions: by comparing using single $L2$ loss, $L2$ with original pairwise learning ($L2+PRL$) and our AdaPRL method, the AdaPRL achieved the highest accuracy among all other methods in Movielens-1M. Specifically, the test MSE and MAE metrics were reduced by 14.82% and 4.74%, 2.05% and 1.43%, and 0.46% and 0.34% respectively with $L1$, $Huber$ and $L2$. AdaPRL also achieved the best performance in the KuaiRec dataset when combined with three different $L_{reg}$. The test MSE and MAE metrics were reduced by 5.01% and 6.76%, 1.51% and 2.42%, and 0.83% and 6.93% respectively with $L1$, $Huber$ and $L2$. Moreover, the Kendall’s $\tau$ was increased by 0.45% and 3.05% respectively, on the ML-1M and KuaiRec datasets with $L2$, indicating that the model not only generates more accurate predictions but also provides a stronger monotonic relationship between labels and predictions.
A.2 Time Series Forecasting